A 28.0 kg block is connected to an empty 1.80 kg b
Answer #1

here,

mass of the block , m = 28 kg

mass of the bucket , mb = 1.8 kg

let the mass of sand be ms

coefficient of static friction , us = 0.41

coefficient of kinetic friction , uk = 0.31

(a)

for the block just to move

( mb + ms) *g - us*m *g = 0

( ms + 1.8) *9.8 = 0.41*28* 9.8

ms = 9.78 kg

the mass of sand added is 9.78 kg


(b)

the accelration of system , a = ( ( ms + mb)*g - uk * m*g) /( m + mb + ms)

a = ( ( 9.78 + 1.8 ) *9.8 - 0.31*28 * 9.8) /( 28 + 9.78 + 1.8)

a = 0.72 m/s^2

the accelration of the system is 0.72 m/s^2